# LongForm: Effective Instruction Tuning with Reverse Instructions

### v2

@@ Appendix B Multilingual Prediction
Appendix C Smaller LongForm Models
@@
Appendix D Qualitative Examples
@@
In a pot, add a can of mushroom soup and 2 cups of beef broth. Heat to a simmer and add 3 cups of sliced mushrooms. Let simmer for 15 minutes. Serve over cooked noodles or rice. Enjoy!
Put the meat in a large saucepan or skillet and cover it with water. Add 1 tsp (5 ml) of salt and bring it to a boil. Remove the meat from the pan and discard the water. Add 1/2 can (125 ml) of mushroom soup and 2 cups (473 ml) of mushrooms and simmer it for 15 minutes. Sprinkle with salt and pepper and garnish with sour cream and parsley before serving.
@@
Digital cameras capture and transfer the light reflected from an object into a.jpg file using a process called image compression. The image compression algorithm takes a series of pixels from the image and arranges them into blocks of similar colors. It then reduces the amount of data needed to store the image by removing redundant information from the blocks. The final result is an.jpg file that can be stored, edited, and shared.
It doesn’t matter how the camera captures the picture; what it captures is a signal, which is an analogue value. That signal is either converted to digital form by a linear analog-to-digital converter (ADC), usually at the image sensor itself, or it’s first converted to a voltage which is then read in by a linear ADC, either at the image sensor or in the digital part of the image processing pipeline. The digital value of the ADC is converted by software into a.jpg. There’s a lot of clever maths going on behind the scenes; but the crucial part is that it is a digital value which is converted to a.jpg.
@@
It was a momentous occasion, a day the heroes would never forget. They had banded together to save the world from the evil overlord, and their triumph was complete. After the celebrations had died down, however, the heroes realized they had no idea how to run a government. The world needed a new leader, but who should it be? What decisions should they make? The heroes knew they were up to the task, but they needed to think carefully if they wanted to do it right. (85 words)
It was over. The final battle, the last stand, and the victory celebration had concluded and the final tally was in: the villains were dead, their minions were scattered or captured, the last of the evil artifacts were destroyed, and the world had been saved. The cheering heroes took a break from their celebrations to hear the good news from the new king of the newly formed world government. He had good news: the world was once again at peace, all villain threats were eliminated, and a new era of prosperity was just beginning. All that remained were some minor issues of governance and organization. He was currently working on those and would be happy to hand over full control to the new rulers as soon as he finished. The heroes looked at each other with doubt and dismay. They were heroes, not politicians. They couldn’t run a world. They were terrible at that sort of thing. Then the realization struck. “You want us to be your government?” The new king was taken aback. “Yes. That’s why you’re here. To take over for me. All I need is your names and titles and I can hand over all power.” The heroes had a serious crisis. (205 words)
@@
Appendix E Details of LongForm-C
E.1 LLM Query
@@
The template for the instruction style: Instruction: X Output: <corpus_example> What kind of instruction could this be the answer to? X:
The template for the informal chatbot style: You are a chatbot. A user sent you an informal message and your reply is as follows. Message: X Reply: <corpus_example> What is the informal message X? X:
The template for the search engine/query style: You are a search engine. A person queried something in detail and the most relevant document about the query is as follows. Query: X Document: <corpus_example> What is the detailed query X? X:
E.2 WikiHow Templates
@@ E.4 BEA-GEC Templates
Appendix F Manual Analysis of the LongForm-C Dataset
@@
Describe a new research study on brain biometrics.
Credit: Chen Song, University at Buffalo. BUFFALO, N.Y. — To overcome password fatigue, many smartphones include facial recognition, fingerprint scans and other biometric systems. “To the best of our knowledge, this is the first in-depth research study on a truly cancelable brain biometric system. We refer to this as ‘hard cancellation,’ meaning the original brain password can be reset without divulging the user’s identity,” says collaborator Zhanpeng Jin, PhD, associate professor of computer science and engineering at UB. Their work is described in a study that will be presented June 11 at MobiSys 2018, a flagship mobile computing conference hosted in Germany by the Association for Computing Machinery. For a copy of the study, please email Cory Nealon at cmnealon@buffalo.edu. The research was supported by the National Science Foundation (NSF) and NSF’s Center for Identification Technology Research.
@@
Can you tell me if Scotland exists?
"Scotland: it certainly does exist. I can say that safely. That mangling of ’Ypres’ ↑ does nothing for my credibility amongst French-speakers though. Tagged with BlackMill Games, M2H, Verdun.”
@@
What does Disney’s streaming video service mean for Netflix?
Could the Mouses’s gauntlet leave Netflix bruised? That’s what investors are wondering after Disney announced it was not only going to launch its own streaming video service, but also no longer license its popular content to Netflix. The Disney-Netflix deal is slated to end in 2019. At that time, Disney plans to have its own subscription-based service that could cause Netflix subscribers to ditch their \$9.99 monthly fee. And according to a new survey from Fluent, as many as 1 in 5 of U.S.-based subscribers could leave the streaming behemoth. The survey shows that many Netflix users would at least consider cancelling their subscription if Disney pulls its content. According to the survey of 2,893 US adults aged 18 and up, 19 percent of total Netflix subscribers would stop their Netflix subscription if they lose access to Disney’s movies and content; 15 percent of non-parents said they would cancel their subscription; 23 percent of parents of kids ages 0-18 would cancel. What does all of this mean for investors? Potentially a devastating loss in revenue for Netflix. On average, Netflix subscribers pay \$10 per month. If 19 percent of the estimated 50.85 million U.S. subscribers cancel their service, Netflix could lose about \$96.6 million in monthly revenue, or about \$1.2 billion for the year. If Disney proves to be successful, consumers could expect other studios to follow suit, which could either result in Netflix losing subscribers, or those studios struggling to attract new subscribers. According to the survey, 62 percent of Americans aren’t likely to subscribe to an additional service, with cost likely being the driving factor. Still, 4 in 10 Americans are willing to pay for additional streaming services. Of course, Netflix isn’t just going to lie down and let this happen. The streaming giant knows content is king, and has been spending billions of dollars developing its own original series that attract subscribers. The company has nearly \$16 billion earmarked to be spent on content deals in the coming years, and it’s still adding subscribers at an impressive rate, recently exceeding 100 million global subscribers. One question that remains unclear is what happens to the Marvel content. Netflix is in “active discussions” with Disney about keeping Marvel and Star Wars films after 2019 when Disney and Pixar movies won’t be in the Netflix catalog, according to a Reuters report. Andrew Dodson is a journalist who has worked in Michigan and Colorado. Email him at [email protected] Disclosure: Streaming Observer is supported by readers. Articles may contain referral links. For more information, see the disclosure at the bottom of the page.
@@
Appendix G Qualitative Examples

### v3

@@ Appendix C Multilingual News Generation
Appendix D Smaller LongForm Models
@@
Appendix E Qualitative Examples
@@
In a pot, add a can of mushroom soup and 2 cups of beef broth. Heat to a simmer and add 3 cups of sliced mushrooms. Let simmer for 15 minutes. Serve over cooked noodles or rice. Enjoy! Put the meat in a large saucepan or skillet and cover it with water. Add 1 tsp (5 ml) of salt and bring it to a boil. Remove the meat from the pan and discard the water. Add 1/2 can (125 ml) of mushroom soup and 2 cups (473 ml) of mushrooms and simmer it for 15 minutes. Sprinkle with salt and pepper and garnish with sour cream and parsley before serving.
@@
Digital cameras capture and transfer the light reflected from an object into a.jpg file using a process called image compression. The image compression algorithm takes a series of pixels from the image and arranges them into blocks of similar colors. It then reduces the amount of data needed to store the image by removing redundant information from the blocks. The final result is an.jpg file that can be stored, edited, and shared. It doesn’t matter how the camera captures the picture; what it captures is a signal, which is an analogue value. That signal is either converted to digital form by a linear analog-to-digital converter (ADC), usually at the image sensor itself, or it’s first converted to a voltage which is then read in by a linear ADC, either at the image sensor or in the digital part of the image processing pipeline. The digital value of the ADC is converted by software into a.jpg. There’s a lot of clever maths going on behind the scenes; but the crucial part is that it is a digital value which is converted to a.jpg.
@@
It was a momentous occasion, a day the heroes would never forget. They had banded together to save the world from the evil overlord, and their triumph was complete. After the celebrations had died down, however, the heroes realized they had no idea how to run a government. The world needed a new leader, but who should it be? What decisions should they make? The heroes knew they were up to the task, but they needed to think carefully if they wanted to do it right. (85 words) It was over. The final battle, the last stand, and the victory celebration had concluded and the final tally was in: the villains were dead, their minions were scattered or captured, the last of the evil artifacts were destroyed, and the world had been saved. The cheering heroes took a break from their celebrations to hear the good news from the new king of the newly formed world government. He had good news: the world was once again at peace, all villain threats were eliminated, and a new era of prosperity was just beginning. All that remained were some minor issues of governance and organization. He was currently working on those and would be happy to hand over full control to the new rulers as soon as he finished. The heroes looked at each other with doubt and dismay. They were heroes, not politicians. They couldn’t run a world. They were terrible at that sort of thing. Then the realization struck. “You want us to be your government?” The new king was taken aback. “Yes. That’s why you’re here. To take over for me. All I need is your names and titles and I can hand over all power.” The heroes had a serious crisis. (205 words)
@@
Appendix F Details of LongForm-C
F.1 LLM Query
@@
The template for the instruction style: {mdframed}[hidealllines=true,backgroundcolor=gray!5] Instruction: X
Output: <corpus_example>
What kind of instruction could this be the answer to?
X:
The template for the informal chatbot style: {mdframed}[hidealllines=true,backgroundcolor=gray!5] You are a chatbot. A user sent you an informal message and your reply is as follows.
Message: X
Reply: <corpus_example>
What is the informal message X?
X:
The template for the search engine/query style: {mdframed}[hidealllines=true,backgroundcolor=gray!5] You are a search engine. A person queried something in detail and the most relevant document about the query is as follows.
Query: X
Document: <corpus_example>
What is the detailed query X?
X:
F.2 WikiHow Templates
@@ F.4 BEA-GEC Templates
Appendix G Diversity Analysis: LongForm-C, Dolly, FLAN
To further analyze the diversity of instructions in LongForm-C, we compare its distribution of noun+verb and auxiliary+verb pairs (Figure 6(a)) with those of the Dolly human annotation dataset (Figure 6(b)) and the FLAN dataset (Figure 6(c)), which focuses specifically on NLP tasks. Figure 6 presents these three distributions side-by-side for easy comparison.
(a) LongForm-C
(b) Dolly
(c) FLAN
Figure 6: Comparison of instruction diversity across LongForm-C, Dolly, and FLAN datasets. Each subfigure shows the distribution of noun+verb and auxiliary+verb pairs extracted from instructions.
We observe that LongForm-C exhibits a variety of instruction types, covering areas such as description, explanation, analysis, and creative tasks. In contrast, the Dolly dataset appears more focused on listing, naming, extraction, and identifying tasks, while still maintaining diversity. The FLAN dataset, being NLP-specific, shows a concentration around text classification, question answering, and sentiment analysis tasks.
This comparison highlights how reverse instructions in LongForm-C generate diverse tasks that complement existing datasets like Dolly and FLAN, suggesting potential benefits in combining multiple data collection strategies for more comprehensive coverage.
Appendix H Manual Analysis of the LongForm-C Dataset
@@
Describe a new research study on brain biometrics. Credit: Chen Song, University at Buffalo. BUFFALO, N.Y. — To overcome password fatigue, many smartphones include facial recognition, fingerprint scans and other biometric systems. “To the best of our knowledge, this is the first in-depth research study on a truly cancelable brain biometric system. We refer to this as ‘hard cancellation,’ meaning the original brain password can be reset without divulging the user’s identity,” says collaborator Zhanpeng Jin, PhD, associate professor of computer science and engineering at UB. Their work is described in a study that will be presented June 11 at MobiSys 2018, a flagship mobile computing conference hosted in Germany by the Association for Computing Machinery. For a copy of the study, please email Cory Nealon at cmnealon@buffalo.edu. The research was supported by the National Science Foundation (NSF) and NSF’s Center for Identification Technology Research.
@@
Can you tell me if Scotland exists? "Scotland: it certainly does exist. I can say that safely. That mangling of ’Ypres’ ↑ does nothing for my credibility amongst French-speakers though. Tagged with BlackMill Games, M2H, Verdun.”
@@
What does Disney’s streaming video service mean for Netflix? Could the Mouses’s gauntlet leave Netflix bruised? That’s what investors are wondering after Disney announced it was not only going to launch its own streaming video service, but also no longer license its popular content to Netflix. The Disney-Netflix deal is slated to end in 2019. At that time, Disney plans to have its own subscription-based service that could cause Netflix subscribers to ditch their \$9.99 monthly fee. And according to a new survey from Fluent, as many as 1 in 5 of U.S.-based subscribers could leave the streaming behemoth. The survey shows that many Netflix users would at least consider cancelling their subscription if Disney pulls its content. According to the survey of 2,893 US adults aged 18 and up, 19 percent of total Netflix subscribers would stop their Netflix subscription if they lose access to Disney’s movies and content; 15 percent of non-parents said they would cancel their subscription; 23 percent of parents of kids ages 0-18 would cancel. What does all of this mean for investors? Potentially a devastating loss in revenue for Netflix. On average, Netflix subscribers pay \$10 per month. If 19 percent of the estimated 50.85 million U.S. subscribers cancel their service, Netflix could lose about \$96.6 million in monthly revenue, or about \$1.2 billion for the year. If Disney proves to be successful, consumers could expect other studios to follow suit, which could either result in Netflix losing subscribers, or those studios struggling to attract new subscribers. According to the survey, 62 percent of Americans aren’t likely to subscribe to an additional service, with cost likely being the driving factor. Still, 4 in 10 Americans are willing to pay for additional streaming services. Of course, Netflix isn’t just going to lie down and let this happen. The streaming giant knows content is king, and has been spending billions of dollars developing its own original series that attract subscribers. The company has nearly \$16 billion earmarked to be spent on content deals in the coming years, and it’s still adding subscribers at an impressive rate, recently exceeding 100 million global subscribers. One question that remains unclear is what happens to the Marvel content. Netflix is in “active discussions” with Disney about keeping Marvel and Star Wars films after 2019 when Disney and Pixar movies won’t be in the Netflix catalog, according to a Reuters report. Andrew Dodson is a journalist who has worked in Michigan and Colorado. Email him at [email protected] Disclosure: Streaming Observer is supported by readers. Articles may contain referral links. For more information, see the disclosure at the bottom of the page.
@@
Appendix I Qualitative Examples